OpenStudy (anonymous):

For each family of functions that depends on one or more parameters, determine the function’s absolute maximum and absolute minimum on the given interval. b) w(x) = a(1- e^-bx), [b , 3b]

OpenStudy (lastdaywork):

For \[x \ge 0\] \[e^x \ge 1\] Now, can you solve the question ??

OpenStudy (anonymous):

I don't get it. Im really not good at math :(

OpenStudy (anonymous):

so x>1 means x= b is max?

OpenStudy (lastdaywork):

How do you generally solve such questions ??

OpenStudy (anonymous):

I still don't know

OpenStudy (anonymous):

my instructor doesn't teach, it is a group work. we have to study by our own.

OpenStudy (lastdaywork):

Can you make a graph for - \[w(x) = 1 - \frac{ 1 }{ e^{bx} } ;\] b/w x = b and x = 3b for case (1) b > 0 (2) b <0 ??

OpenStudy (anonymous):

b>0 min and for b<0 is max? is that right?

OpenStudy (lastdaywork):

You'll get the same answer for both cases.

OpenStudy (anonymous):

so its neither

OpenStudy (lastdaywork):

IDK how to explain you without referring to the graph. Maybe @agent0smith can help here.